# Mean-field limits for Consensus-Based Optimization and Sampling

### v2

@@ 2.5 Quantitative mean-field limit for CBS
Using the same reasoning as that in the proof of Theorem 2.6, we can prove a mean-field result for the CBS dynamics (1.4). We state the theorem and, given the similarity of the proof with that of Theorem 2.6, only list the key ingredients needed for the proof.
@@
In the same way as for CBO, it is possible to prove the following sharper statement. Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$. Consider the system (2.8) and (2.9) with the coefficients given by (2.18). If $q\geqslant 4p_{\mathcal{M}}(s,\ell)$, then for all $p\in(0,\frac{q}{2}]$, there is $C>0$ independent of $J$ such that
@@
  • Wasserstein stability estimate for the weighted covariance. We prove in Corollary 3.3 that, for all $R>0$ and for all $p\geqslant 2p_{\mathcal{M}}(\ell,s)$, there exists $\widetilde{L}_{\sqrt{\mathcal{C}}}(R,p)>0$ such that
@@ 3 Auxiliary results
This section is organized as follows. In Subsection 3.1, we improve on existing Wasserstein stability estimates for the weighted mean and prove novel ones for the weighted covariance and the square root of the weighted covariance. In Subsection 3.2, we present moment estimates for the empirical measures associated with the CBO (1.1) and CBS (1.4) dynamics. In Subsection 3.3, we prove convergence estimates for the weighted moments when particle positions are drawn independently from a given probability distribution, using results from the statistics literature. Finally, in Subsection 3.4, we prove that if $\mathcal{C}_{\beta}(\overline{\rho}_{0})\succ 0$, then the covariance matrix $\mathcal{C}_{\beta}(\overline{\rho}_{t})\succ 0$ associated with the solution to the mean-field CBS dynamics (1.5) remains strictly positive definite for all times.
@@ 3.1 Wasserstein stability estimates for the weighted moments
In particular $\nu_{i}[\lvert x\rvert^{p}]\to\infty$ in the limit as $i\to\infty$, otherwise the left-hand side of 3.11 would be bounded from above uniformly in $i$ by (3.8). Therefore $W_{p}(\nu_{i},\delta_{0})\to\infty$. On the other hand, by the triangle inequality,
@@ 3.2 Moment estimates for the empirical measures
By assumption it holds that $\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}\succcurlyeq\sqrt{\eta}\operatorname{I}_{d}$. By an inequality due to van Hemmen and Ando [vanHemmen1980], see also [MR1477662, Problem X.5.5], it holds that
@@ 3.4 No collapse in finite time for CBS
In order to deal with the matrix square root in the CBS dynamics, we prove that the weighted covariance stays positive definite for finite times.
@@
In particular, if $\mathcal{C}(\overline{\rho}_{0})\succ 0$ and if $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ is continuous, then
@@
Thus we obtain $y^{\mathsf{T}}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\bigr{)}=0$ for $\overline{\rho}_{t}$-almost every $x\in\mathbf{R}^{d}$, and so $y^{\mathsf{T}}\mathcal{M}_{\beta}(\overline{\rho}_{t})=y^{\mathsf{T}}\mathcal{M}(\overline{\rho}_{t})$ by integrating against $\overline{\rho}_{t}$. But then,
$y^{\mathsf{T}}\mathcal{C}(\overline{\rho}_{t})y=\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\,\overline{\rho}_{t}(\mathrm{d}x)=\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\,\overline{\rho}_{t}(\mathrm{d}x)=0.$
which is a contradiction. Thus $C_{\beta}(\overline{\rho}_{t})\succ 0$ for all $t\in[0,T]$, which implies the second inequality in (3.14). ∎
@@ 5 Extensions and perspectives for future work
This section is organized as follows. First, in Subsection 5.1, we present an alternative approach based on stopping times for proving our main results Theorems 2.6 and 2.12. Then, in Subsection 5.2, we show how the stopping time approach presented in Subsection 5.1 can be employed in order to obtain a mean-field result in the setting where $f\in\mathcal{A}(s,\ell,u)$ with $u>\ell>0$, under which the improved stability estimates in Corollary 3.3 are no longer valid.
@@ 5.1 Proving mean-field results using stopping times
We illustrate the stopping times approach only for Theorem 2.12, noting that Theorem 2.6 can be proved in exactly the same manner.
@@
Let $R>0$ be such that the stopping time
$\theta_{J}(R):=\inf\Bigl{\{}t\geqslant 0:\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}\right\rvert^{p}\geqslant R\Bigr{\}},$ (5.2)
satisfies $\mathbf{P}[\theta_{J}\leqslant T]=\mathcal{O}(J^{-2}).$ Note that the event $\{\theta_{J}\leqslant T\}$ coincides with the event $\Omega_{J}$ given in (5.1). From Hölder’s inequality it follows that
@@
$\left\lvert X_{t\wedge\theta_{J}}^{(j)}-\overline{X}_{t\wedge\theta_{J}}^{(j)}\right\rvert^{p}\leqslant$ $\left\lvert\int_{0}^{t\wedge\theta_{J}}b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}s\right\rvert^{p}$
$+\left\lvert\int_{0}^{t\wedge\theta_{J}}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}\right\rvert^{p}.$
@@
$M_{t}=\int_{0}^{t}\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\,\mathrm{d}W_{s}.$
@@
$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant(2T)^{p-1}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$
$\qquad+C_{\rm BDG}2^{p-1}T^{\frac{p}{2}-1}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (5.4)
From the triangle inequality we obtain that
$\frac{1}{2^{p-1}}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$ $\leqslant\int_{0}^{t}\mathbf{E}\left\lvert b\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-b\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rvert^{p}\,\mathrm{d}s$
$\qquad+\int_{0}^{t}\mathbf{E}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s.$ (5.5)
@@
$\frac{1}{2^{p-1}}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s$ $\leqslant\int_{0}^{t}\mathbf{E}\left\lVert\sigma\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-\sigma\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s$
$\qquad+\int_{0}^{t}\mathbf{E}\left\lVert\sigma\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (5.6)
@@
    $\left\lVert\sigma\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-\sigma\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rVert_{\rm F}^{p}$ $=\left(2\lambda^{-1}\right)^{\frac{p}{2}}\biggl{\lVert}\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{s\wedge\theta_{J}}\right)}-\sqrt{\mathcal{C}_{\beta}\left(\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)}\biggr{\rVert}_{\rm F}^{p}.$
@@
    $\mathbf{E}\left[W_{p}\left(\mu^{J}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\right\rvert^{p},$
@@
    $\mathbf{E}\left\lVert\sigma\left(X_{s\wedge\theta_{J}}^{(j)},\mu^{J}_{s\wedge\theta_{J}}\right)-\sigma\left(\overline{X}_{s\wedge\theta_{J}}^{(j)},\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)\right\rVert_{\rm F}^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p}.$
  • For the other diffusion term, we use Lemma 3.8. The assumptions of this lemma are satisfied since, on the one hand, $\overline{\rho}_{s}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ by Theorem 2.4 and $2p<q$, and on the other hand, it holds that $\mathcal{C}_{\beta}(\overline{\rho}_{s})\succ 0$ by virtue of Proposition 3.9. We deduce that
    $\mathbf{E}\left\lVert\sigma\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}=\left(2\lambda^{-1}\right)^{\frac{p}{2}}\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}\left(\overline{\mu}^{J}_{s}\right)}-\sqrt{\mathcal{C}_{\beta}\left(\overline{\rho}_{s}\right)}\right\rVert_{\rm F}^{p}\leqslant CJ^{-\frac{p}{2}}.$
@@
$\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s\leqslant CJ^{-\frac{p}{2}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u\wedge\theta_{J}}-\overline{X}^{(j)}_{u\wedge\theta_{J}}\right\rvert^{p}\right]\,\mathrm{d}s.$
Applying a similar reasoning to the drift terms, we finally obtain from (5.4) that
$\mathbf{E}\left[\sup_{s\in[0,t]}\Big{\lvert}X_{s\wedge\theta_{J}}^{(j)}-\overline{X}_{s\wedge\theta_{J}}^{(j)}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\bigl{\lvert}X_{u\wedge\theta_{J}}^{(j)}-\overline{X}_{u\wedge\theta_{J}}^{(j)}\bigr{\rvert}^{p}\right]\,\mathrm{d}s.$
@@
In addition, by Lemma 2.5 it holds that
$\mathbf{P}[\theta_{J}(R)\leqslant T]=\mathbf{P}\left[\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}\geqslant R\right]\leqslant CJ^{-2}.$

### v3

@@ 2.5 Quantitative mean-field limit for CBS
Proving a mean-field limit for the CBS dynamics (1.4) is more challenging than for CBO because of the square root of the weighted covariance in the diffusion coefficient. Thanks to our novel stability estimate for the square root of the weighted covariance (see Corollary 3.3 below), we can prove a mean-field result for the CBS dynamics (1.4) in the same way as for CBO.
@@
In the same way as for CBO, it is possible to prove the following more general statement. Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$. If $q\geqslant 4p_{\mathcal{M}}(s,\ell)$, then for all $p\in(0,\frac{q}{2}]$, there is $C>0$ independent of $J$ such that
@@
  • Wasserstein stability estimate for the weighted covariance. The crucial ingredient for the proof of Theorem 2.10 is a stability estimate for the square root of the weighted covariance. We prove in Corollary 3.3 that, for all $R>0$ and for all $p\geqslant 2p_{\mathcal{M}}(\ell,s)$, there exists $\widetilde{L}_{\sqrt{\mathcal{C}}}(R,p)>0$ such that
@@
With these ingredients at hand, we can prove Theorem 2.10 in the same way as for CBO, so we skip the proof here. Let us mention, however, that a different proof based on a technique using stopping times is presented in Subsection 5.1.
@@ 3 Auxiliary results
This section is organized as follows. In Subsection 3.1, we improve on existing Wasserstein stability estimates for the weighted mean and prove novel ones for the weighted covariance and the square root of the weighted covariance. Then, in Subsection 3.2, we present moment estimates for the empirical measures associated with the CBO (1.1) and CBS (1.4) dynamics. In Subsection 3.3, we prove convergence estimates for the weighted moments when particle positions are drawn independently from a given probability distribution, using results from the statistics literature. Finally, in Subsection 3.4, we prove that if $\mathcal{C}_{\beta}(\overline{\rho}_{0})\succ 0$, then the covariance matrix $\mathcal{C}_{\beta}(\overline{\rho}_{t})\succ 0$ associated with the solution to the mean-field CBS dynamics (1.5) remains strictly positive definite for all times.
@@ 3.1 Wasserstein stability estimates for the weighted moments
In particular $\nu_{i}[\lvert x\rvert^{p}]\to\infty$ in the limit as $i\to\infty$, otherwise the left-hand side of 3.11 would be bounded from above uniformly in $i$ by (3.8). Therefore, $W_{p}(\nu_{i},\delta_{0})\to\infty$. On the other hand, by the triangle inequality,
@@ 3.2 Moment estimates for the empirical measures
By assumption, it holds that $\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}\succcurlyeq\sqrt{\eta}\operatorname{I}_{d}$. By an inequality due to van Hemmen and Ando [vanHemmen1980], see also [MR1477662, Problem X.5.5], it holds that
@@ 3.4 No collapse in finite time for CBS
In order to deal with the matrix square root in the CBS dynamics, we prove that the weighted covariance stays positive definite for finite times. This is particularly relevant when CBS is run in optimization mode with $\lambda=1$, since then $\mathcal{C}(\overline{\rho}_{t})\to 0$ as $t\to\infty$ under appropriate assumptions [CBS-Carrillo2021].
@@
In particular, if $\mathcal{C}(\overline{\rho}_{0})\succ 0$ and if $t\mapsto\mathcal{C}_{\beta}(\rho_{t})$ is continuous (which is true under the assumptions of Theorem 2.4), then
@@
Thus, we obtain $y^{\mathsf{T}}\bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\bigr{)}=0$ for $\overline{\rho}_{t}$-almost every $x\in\mathbf{R}^{d}$, and so $y^{\mathsf{T}}\mathcal{M}_{\beta}(\overline{\rho}_{t})=y^{\mathsf{T}}\mathcal{M}(\overline{\rho}_{t})$ by integrating against $\overline{\rho}_{t}$. But then,
$y^{\mathsf{T}}\mathcal{C}(\overline{\rho}_{t})y=\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\,\overline{\rho}_{t}(\mathrm{d}x)=\int_{\mathbf{R}^{d}}\left\lvert y^{\mathsf{T}}\Bigl{(}x-\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{)}\right\rvert^{2}\,\overline{\rho}_{t}(\mathrm{d}x)=0,$
which is a contradiction. Thus, $C_{\beta}(\overline{\rho}_{t})\succ 0$ for all $t\in[0,T]$, which implies the second inequality in (3.14). ∎
@@ 5 Extensions and perspectives for future work
This section is organized as follows. First, in Subsection 5.1, we present an alternative approach based on stopping times for proving our main results Theorems 2.6 and 2.10. Then, in Subsection 5.2, we show how the stopping time approach presented in Subsection 5.1 can be employed in order to obtain a mean-field result in the setting where $f\in\mathcal{A}(s,\ell,u)$ with $u>\ell>0$, in which the improved stability estimates in Corollary 3.3 are no longer valid.
@@ 5.1 Proving mean-field results using stopping times
We illustrate the stopping times approach only for Theorem 2.10, noting that Theorem 2.6 can be proved in exactly the same manner. Furthermore, we assume for simplicity that $\overline{\rho}_{0}$ has infinitely many moments.
@@
For $R>0$, let $\theta_{J}(R)$ be the stopping time
$\theta_{J}(R):=\inf\Bigl{\{}t\geqslant 0:\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}\geqslant R\Bigr{\}},$ (5.2)
Note that the event $\{\theta_{J}\leqslant T\}$ coincides with the event $\Omega_{J}$ given in (5.1). By Hölder’s inequality, it holds for all $q>p$ that
@@
$\frac{1}{2^{p-1}}\left\lvert X_{t\wedge\theta_{J}}^{(j)}-\overline{X}_{t\wedge\theta_{J}}^{(j)}\right\rvert^{p}\leqslant$ $\left\lvert\int_{0}^{t\wedge\theta_{J}}b\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}s\right\rvert^{p}$
$+\left\lvert\int_{0}^{t\wedge\theta_{J}}\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}W_{s}\right\rvert^{p}.$
@@
$M_{t}=\int_{0}^{t}\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\,\mathrm{d}W_{s}.$
@@
$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\bigr{\rvert}^{p}\right]$ $\leqslant(2T)^{p-1}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lvert b\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s$
$\qquad+C_{\rm BDG}2^{p-1}T^{\frac{p}{2}-1}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (5.4)
Applying the triangle inequality gives that
$\frac{1}{2^{p-1}}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\bigl{\lvert}b\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\bigr{\rvert}^{p}\,\mathrm{d}s$ $\leqslant\int_{0}^{t}\mathbf{E}\left\lvert b\Bigl{(}X^{(j)}_{s\wedge\theta_{J}},\mu^{J}_{s\wedge\theta_{J}}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s$
$\qquad+\int_{0}^{t}\mathbf{E}\left\lvert b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\mu}^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s.$ (5.5)
@@
$\frac{1}{2^{p-1}}\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s$ $\leqslant\int_{0}^{t}\mathbf{E}\left\lVert\sigma\Bigl{(}X^{(j)}_{s\wedge\theta_{J}},\mu^{J}_{s\wedge\theta_{J}}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s$
$\qquad+\int_{0}^{t}\mathbf{E}\left\lVert\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\mu}^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (5.6)
@@
    $\left\lVert\sigma\Bigl{(}X^{(j)}_{s\wedge\theta_{J}},\mu^{J}_{s\wedge\theta_{J}}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\Bigr{)}\right\rVert_{\rm F}^{p}$ $=\left(2\lambda^{-1}\right)^{\frac{p}{2}}\biggl{\lVert}\sqrt{\mathcal{C}_{\beta}\left(\mu^{J}_{s\wedge\theta_{J}}\right)}-\sqrt{\mathcal{C}_{\beta}\left(\overline{\mu}^{J}_{s\wedge\theta_{J}}\right)}\biggr{\rVert}_{\rm F}^{p}.$
@@
    $\mathbf{E}\Bigl{[}W_{p}\bigl{(}\mu^{J}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\bigr{)}^{p}\Bigr{]}\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{k=1}^{J}\left\lvert X^{(k)}_{s\wedge\theta_{J}}-\overline{X}^{(k)}_{s\wedge\theta_{J}}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p},$
@@
    $\mathbf{E}\left\lVert\sigma\Bigl{(}X^{(j)}_{s\wedge\theta_{J}},\mu^{J}_{s\wedge\theta_{J}}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s\wedge\theta_{J}},\overline{\mu}^{J}_{s\wedge\theta_{J}}\Bigr{)}\right\rVert_{\rm F}^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\right\rvert^{p}.$
  • For the other diffusion term, we use Lemma 3.8. The assumptions of this lemma are satisfied, since it holds that $\overline{\rho}_{s}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ for all $q>0$ by Theorem 2.4, and additionally $\mathcal{C}_{\beta}(\overline{\rho}_{s})\succ 0$ by Proposition 3.9. We deduce that
    $\mathbf{E}\left\lVert\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\mu}^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}=\left(2\lambda^{-1}\right)^{\frac{p}{2}}\mathbf{E}\left\lVert\sqrt{\mathcal{C}_{\beta}\left(\overline{\mu}^{J}_{s}\right)}-\sqrt{\mathcal{C}_{\beta}\left(\overline{\rho}_{s}\right)}\right\rVert_{\rm F}^{p}\leqslant CJ^{-\frac{p}{2}}.$
@@
$\mathbf{E}\int_{0}^{t\wedge\theta_{J}}\left\lVert\sigma\Bigl{(}X^{(j)}_{s},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}^{(j)}_{s},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s\leqslant CJ^{-\frac{p}{2}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u\wedge\theta_{J}}-\overline{X}^{(j)}_{u\wedge\theta_{J}}\right\rvert^{p}\right]\,\mathrm{d}s.$
Applying a similar reasoning for the drift terms, we finally obtain from (5.4) that
$\mathbf{E}\left[\sup_{s\in[0,t]}\Big{\lvert}X^{(j)}_{s\wedge\theta_{J}}-\overline{X}^{(j)}_{s\wedge\theta_{J}}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\Bigl{\lvert}X_{u\wedge\theta_{J}}^{(j)}-\overline{X}_{u\wedge\theta_{J}}^{(j)}\Bigr{\rvert}^{p}\right]\,\mathrm{d}s.$
@@
In addition, by Lemma 2.5, it holds for any sufficiently large fixed $R$, that for all $\alpha>0$ there is $C_{\alpha}$ such that
$\mathbf{P}[\theta_{J}(R)\leqslant T]=\mathbf{P}\left[\sup_{t\in[0,T]}\frac{1}{J}\sum_{j=1}^{J}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}\geqslant R\right]\leqslant C_{\alpha}J^{-\alpha}.$